Natalie is thinking of buying a van that will be used only for business. The cost of the van is estimated at $36,500. Natalie would spend an additional $2,500 to have the van painted. In addition, she wants the back seat of the van removed so that she will have lots of room to transport her mixer inventory as well as her baking supplies. The cost of taking out the back seat and installing shelving units is estimated at $1,500. She expects the van to last about 5 years, and she expects to drive it for 200,000 miles. The annual cost of vehicle insurance will be $2,400. Natalie estimates that at the end of the 5-year useful life the van will sell for $7,500. Assume that she will buy the van on August 15, 2016, and it will be ready for use on September 1, 2016.

Natalie is concerned about the impact of the van’s cost on her income statement and balance sheet. She has come to you for advice on calculating the van’s depreciation.

Instructions

(a) Determine the cost of the van.

(b) Prepare three depreciation tables for 2016, 2017, and 2018: one for straight-line depreciation (similar to the one in Illustration 9-10), one for double-declining balance depreciation (Illustration 9-14), and one for units-of-activity depreciation (Illustration 9-12). For units-of-activity, Natalie estimates she will drive the van as follows: 15,000 miles in 2016; 50,000 miles in 2017; 50,000 miles in 2018; 45,000 miles in 2019; 40,000 miles in 2020. Recall that Cookie Creations has a December 31 year-end.

(c) What impact will the three methods of depreciation have on Natalie’s balance sheet at December 31, 2016? What impact will the three methods have on Natalie’s income statement in 2016?

(d) What impact will the three methods of depreciation have on Natalie’s income statement over the van’s total 5-year useful life?

(e) What method of depreciation would you recommend Natalie use?
Answer #1
a) Cost of Van:
Purchase Cost 36500
Add: Cost of painting 2500
Add: Cost of taking out back seat and installing shelving units 1500
40500
Usefull life 5
Expected drive 200000 mile
annual cost of insurance 2400
Expected Resale value after usefull life 7500
purchase date 15th august 2016
ready for use date (This date is to be used for calculating depreciation) 1st September 2016

b)

SLM Method depreciation Depreciation amount
= (Cost of asset - Resale value)/use full life
per year depreciation 6600
Depreciation for year 2016
for the period 1st september to 31st december 2200
(6600/12*4 months)
Depreciation for year 2017 (For full year) 6600
Depreciation for year 2018 (For full year) 6600
Depreciation for year 2019 (For full year) 6600
Depreciation for year 2020 (For full year) 6600
Depreciation for year 2021 (For january 21 to august 21) 4400
Double declining balance method Formula Depreciation
Normal Depreciation rate = (1/usefull life) * 100
= (1/5)*100 20.00%
double declining depreciation rate = 2 * 20% 40%
Depreciation for year 2016
for the period 1st september to 31st december 2016 = cost of assets/declined value * dep. Rate
= 40500 * 40% * 4/12 5400
Depreciation for year 2017 = (40500-5400) * 40% 14040
Depreciation for year 2018 = (40500-5400 - 14040) * 40% 8424
Depreciation for year 2019 = (40500-5400 - 14040- 8424) * 40% 5054.4
Depreciation for year 2020 = (40500-5400 - 14040- 8424-5054.4) * 40% 3032.64
Depreciation for year 2021 = (40500-5400 - 14040- 8424-5054.4-3032.64) * 40% 1819.584
Unit of activity depreciation Formula Depreciation
per Mile depreciation rate = (cost of assets- salvage value)/exepcted units
= (40500 - 7500)/ 200000 0.165
Depreciation for year 2016 (per unit depreciaton rate * units of year
= (0.165* 15000) 2475
Depreciation for year 2017 = (0.165* 50000) 8250
Depreciation for year 2018 = (0.165* 50000) 8250
Depreciation for year 2019 = (0.165* 45000) 7425
Depreciation for year 2020 = (0.165* 40000) 6600

c)

SLM Method
book value of van in balance sheet as on 31st december 2016 38300
Depreciation on income statement in 2016 2200
Double declining balance method
book value of van in balance sheet as on 31st december 2016 35100
Depreciation on income statement in 2016 5400
Unit of activity depreciation
book value of van in balance sheet as on 31st december 2016 38025
Depreciation on income statement in 2016 2475

d)

SLM Method impact on income statement Amount
Depreciation on income statement in 2016 2200
Depreciation on income statement in 2017 6600
Depreciation on income statement in 2018 6600
Depreciation on income statement in 2019 6600
Depreciation on income statement in 2020 6600
Depreciation on income statement in 2021 4400
Double declining balance method impact on income statement Amount
Depreciation on income statement in 2016 5400
Depreciation on income statement in 2017 14040
Depreciation on income statement in 2018 8424
Depreciation on income statement in 2019 5054.4
Depreciation on income statement in 2020 3032.64
Depreciation on income statement in 2021 1819.584
Unit of activity depreciation impact on income statement Amount
Depreciation on income statement in 2016 2475
Depreciation on income statement in 2017 8250
Depreciation on income statement in 2018 8250
Depreciation on income statement in 2019 7425
Depreciation on income statement in 2020 6600

e)

we would like to suggest unit of activity method of depreciation as it will depreciate all value asset in the year 2020 itself.